### v1

# Risk-Calibrated Affective Speech Recognition via Conformal Coverage Guarantees: A Stochastic Calibrative Framework for Emergent Uncertainty Quantification

Abstract
Traffic safety challenges arising from extreme driver emotions highlight the urgent need for reliable emotion recognition systems. Traditional deep learning approaches in speech emotion recognition suffer from overfitting and poorly calibrated confidence estimates. We propose a framework integrating Conformal Prediction (CP) and Risk Control, using Mel-spectrogram features processed through a pre-trained convolutional neural network. Our key innovation is the development of a nonconformity score that heuristically measures how closely a classifier’s predictions align with given inputs. Through calibration samples, we compute this score and derive a statistically rigorous threshold based on user-specified risk level $\alpha$, constructing prediction sets with provable coverage guarantees ($\geq 1-\alpha$). The Risk Control framework enables task-specific adaptation through customizable loss functions, dynamically adjusting prediction set sizes while maintaining coverage guarantees. Cross-dataset experiments on IEMOCAP and TESS demonstrate: 1) Strict coverage guarantee, 2) Significant negative correlation between Average Prediction Set Size (APSS) and $\alpha$, revealing reduced model uncertainty under high-risk conditions. We further propose APSS as a novel metric for evaluating classification uncertainty. This approach enhances speech emotion recognition reliability, with direct applications in intelligent transportation systems and real-time emotion monitoring.
keywords:
Speech emotion recognition; Conformal Prediction; Risk Control; Mel spectrum; Prediction set coverage guarantee
[…]
[label1] organization=School of Automation Science and Electrical Engineering, Beihang University, addressline=No. 37 Xueyuan Road, Haidian District, city=Beijing, postcode=100191, state=Haidian District, country=China
1 Introduction
In recent years, road rage has become an increasingly serious traffic safety problem worldwide.With the acceleration of urbanization and the increase in car ownership, traffic congestion and driving pressure are increasing, resulting in more and more drivers showing extreme emotional reactions while driving.Studies have shown that road rage not only increases the incidence of traffic accidents, but also exacerbates conflicts between drivers and other road users, further threatening public safety. According to the National Highway Traffic Safety Administration, more than one-third of traffic accidents are related to drivers losing control of their emotions, and this proportion is on the rise.
Speech emotion recognition technology, as a means of emotion recognition based on speech signal analysis, has gradually become an effective tool for solving problems such as road rage. By capturing the driver’s voice characteristics in real time and analyzing his or her emotional state, speech emotion recognition can provide timely emotional feedback to the traffic management system, helping to predict and intervene in potential dangerous driving behaviors. This technology can not only provide more accurate emotion monitoring in smart cars, but can also be applied to other emotion-related fields, such as customer service and mental health monitoring. Therefore, the importance of speech emotion recognition is increasing day by day, becoming one of the core research directions in the fields of intelligent transportation systems, unmanned driving technology, and emotion analysis.
With the rapid development of deep learning technology, speech emotion recognition tasks based on convolutional neural networks have made significant progress. This paper extracts the Mel-spectrogram features of speech signals to characterize different emotion categories and conducts experimental comparisons. Despite this, the model may still encounter overfitting problems during training, resulting in performance degradation on new and unseen data. In addition, current deep learning models generally lack confidence or uncertainty measures for prediction results, which means that in practical applications, the model’s predictions may be unreliable. Recent studies have pointed out that the miscalibration problem still exists. For example, Zhao et al. (2023) pointed out that although convolutional neural networks show high accuracy in speech emotion recognition tasks, the confidence scores of their prediction results are often too high and fail to truly reflect the actual performance of the model, resulting in incorrect emotion recognition. Therefore, improving the uncertainty measure of the model has become the key to improving the performance of speech emotion recognition systems Psaros et al. (2023); Wang et al. (2025b).
To solve these problems, the conformal prediction method provides a new solution. CP can not only give a predicted label, but also build a set of all possible prediction results and provide a confidence guarantee for each prediction result. To ensure that the confidence of the selected label meets the risk tolerance specified by the user, specifically, the Mel-spectrogram feature is used as the feature extraction method in the speech emotion recognition task, and 5 neural networks with basically the same parameter quantity are trained on the IEMOCAP dataset to obtain a pre-trained model. After the feature extraction and model training are completed, the Conformal Prediction framework is used to evaluate the model to ensure the reliability of the prediction results. Specifically, by defining non-conformal scores and associating these scores with the real emotion labels, the calibration set is used to calibrate the model. In this process, the test data is processed according to the risk level (alpha value) specified by the user, and finally a prediction set with marginal coverage guarantee is generated. This process ensures that the emotion recognition model not only gives the emotion label, but also provides a confidence estimate for the label, thereby ensuring the statistical reliability of the prediction results.
Although Conformal Prediction can effectively provide coverage guarantee, it only solves the coverage problem of the prediction set and fails to consider the reliability of the model under different risk levels. In order to solve this problem, the Risk Control framework is further introduced. The core idea of this framework is to adjust the risk level of the model on the basis of maintaining coverage guarantee to cope with various uncertainties that may arise in practical applications. Under the Risk Control framework, the beta value is used to quickly determine the optimal prediction set size through the binary search method. By processing the calibration set, the size of the prediction set can be adjusted under the constraint of a given risk level to meet the needs of risk control. At the same time, the loss function is also improved so that it not only focuses on the prediction error rate, but also considers the size and uncertainty of each prediction set. The improvement of the loss function enables the framework to be flexibly applied to different tasks and scenarios, such as emotion recognition, intelligent traffic monitoring, etc. By combining Conformal Prediction and Risk Control, not only the coverage of the model is guaranteed, but also the robustness and accuracy of the model are improved, which can adapt to different practical application needs. Experimental results show that this framework can maintain high-precision emotion recognition at different risk levels while reducing the overfitting and uncertainty of the model, making it more valuable for practical applications.
In order to verify the effectiveness of the framework, a large number of experiments were conducted. First, starting from the pre-experimental part, the performance of different feature selection methods and pre-trained models was tested. In the pre-experimental stage, a variety of feature extraction methods were compared, including mel-spectrogram features, chromaticity features, and spectral contrast features. Through testing on the TESS dataset, it was found that the performance of the mel-spectrogram feature was better than other feature extraction methods, reaching a test set accuracy of 99.82%. Therefore, the mel-spectrogram feature was finally selected as the speech feature extraction method. Next, several different neural network models, including ResNet50, MobileNetV3, ShuffleNetV2, etc., were trained on the IEMOCAP dataset, and their performance in emotion recognition tasks was tested. Although the accuracy of these models on the training set reached 100%, the accuracy on the test set decreased to 44.75%, 40.56% and 35.06% respectively. This phenomenon shows that the model has a certain overfitting problem when facing unseen data, so the Conformal Prediction framework was introduced to ensure the reliability and statistical significance of the model prediction results. To prove the coverage guarantee of the framework, two rounds of validation were conducted. The first round was conducted on the same dataset to ensure that the model can provide stable performance on the calibration set and the test set, verifying that Conformal Prediction can ensure the accuracy of the prediction results at a given risk level. The second round of experiments conducted cross-dataset validation, that is, the model trained on the IEMOCAP dataset was applied to the TESS dataset to test the generalization ability of the model. The experimental results show that the model can still maintain a high coverage across different datasets and can effectively adapt to different data distributions, verifying the robustness of the framework. Next, the uncertainty of different models is evaluated based on the prediction set size. During the training and testing process of each model, the prediction set size is calculated, and the uncertainty of the model is quantified based on the prediction set size at each alpha value. By conducting multiple experiments under approximately the same parameter conditions, when the uncertainty of the model increases, the prediction set size usually increases. This shows that a larger prediction set means that the model has lower confidence in the recognition of certain emotion categories. In order to control this uncertainty, the Risk Control framework is adopted, combined with an improved loss function to optimize the prediction set size to ensure that the model can provide reliable prediction results at different risk levels. Through these experiments, we not only verified the effectiveness of Conformal Prediction in speech emotion recognition, but also proved the application value of Risk Control framework in reducing uncertainty and improving prediction accuracy. The final experimental results show that the combination of these two frameworks can maintain the high accuracy and robustness of the model under different data sets and risk levels, ensuring that the emotion recognition system can meet the needs of practical applications.
2 Related Work
2.1 Speech Recognition
The history of speech recognition technology can be traced back to the 1950s. The initial systems relied on traditional algorithms based on template matching, which usually required users to pronounce in a specific way. With the improvement of computing power, hidden Markov models (HMMs) have become the mainstream technology for speech recognition, which can effectively model speech signals in time series. However, these traditional methods still face problems with recognition accuracy and noise interference, especially in complex speech environments. The performance of traditional speech recognition systems often cannot meet the requirements of high accuracy and high robustness. After entering the era of deep learning, speech recognition technology based on neural networks has completely changed this situation. In particular, the introduction of convolutional neural networks (CNNs) and recurrent neural networks (RNNs) has made the processing of speech signals more efficient and accurate. Technologies such as deep neural networks (DNNs) and long short-term memory networks (LSTMs) have greatly improved the accuracy and real-time performance of speech recognition by automatically extracting features and performing end-to-end training. Today, deep learning has become the core technology of speech recognition and is widely used in smart assistants, speech-to-text, speech translation and other fields. In terms of speech emotion recognition, the advantages of deep learning are also significant. Traditional emotion recognition methods mostly rely on manually extracted features, such as pitch, volume, rhythm, etc. Although they can provide certain information for emotion classification, they are unable to cope with complex emotions. The emotion recognition system based on deep learning, especially in combination with CNN, LSTM and other network structures, can automatically learn richer emotional features from a large amount of speech data, improving the accuracy and robustness of emotion recognition. This has enabled speech emotion recognition to achieve breakthrough progress in multiple application scenarios, especially in the fields of intelligent transportation, customer service, mental health, etc., showing great application potential.
2.2 Conformal Prediction
Conformal Prediction Angelopoulos and Bates (2021); Ye et al. (2024); Barber et al. (2023); Wang et al. (2024); Sun and Yu (2023); Wang et al. (2025a) is an uncertainty quantification method that provides strict statistical guarantees for machine learning models. Unlike traditional point prediction, CP provides a set of possible prediction labels for each input sample, rather than just a single prediction label, by generating a prediction set containing the prediction results. The core idea is to calculate the nonconformity score to measure the difference between each predicted label and the model’s prediction result, thereby deciding whether the label should be included in the prediction set. The nonconformity score measures the model’s confidence in the prediction result. For classification tasks, it is based on the difference between the model’s predicted probability of the true category and the actual category; for regression tasks, it is the difference between the predicted value and the true value. Before generating the prediction set, Conformal Prediction first uses a calibration set to estimate the distribution of the nonconformity score. The calibration set is a part of the data separated from the training set and is specifically used to measure the prediction uncertainty of the model. By calculating the nonconformity score in the calibration set, Conformal Prediction generates a quantitative value that defines which prediction labels the model should include at a given confidence level. In the test phase, for each new sample, the model outputs a predicted label and its corresponding nonconformity score. Then, Conformal Prediction generates a prediction set by comparing the inconsistency score of each label with the quantized value in the calibration set. The prediction set contains all labels whose inconsistency scores are less than or equal to the quantized value. In this way, the generated prediction set can ensure that the true label always appears in the prediction set at the specified confidence level. This provides a coverage guarantee, that is, at the set confidence level, the probability that the generated prediction set contains the true label is at least the confidence level (such as 90%). This guarantee applies to all models and does not depend on any assumed data distribution, so it is called a distribution-independent prediction method.
3 Method
3.1 Preliminaries
As a statistical consistency prediction method, the SCP (Split Conformal Prediction) framework shows unique advantages in the reliability evaluation of machine learning models by converting heuristic uncertainty measures into prediction sets with strict probability guarantees. Its core idea is based on the assumption of data exchangeability and combines the quantile calibration of nonconformity scores to achieve coverage probability control of unknown samples. The theoretical basis and algorithm flow of SCP will be systematically explained below.
3.1.1 Theoretical Foundations and Algorithm Process
1. Problem Definition and Input Conditions
  • 1.
    Input Data:
    • (a)
      Calibration dataset $\{(X_{t},Y_{t}^{*})\}_{t=1}^{n}$, satisfying the independent and identically distributed (i.i.d.) assumption.
    • (b)
      Pretrained classification model $\hat{f}(\cdot)$, with output as a class probability vector $\hat{f}(X_{t})\in[0,1]^{K}$, where the true label probability is $\hat{f}(X_{t})_{Y_{t}^{*}}$.
  • 2.
    Objective Function: Construct the prediction set $\mathcal{C}(X_{\text{test}})$ such that the coverage probability of the true label satisfies:
    $\mathbb{P}(Y_{\text{test}}^{*}\in\mathcal{C}(X_{\text{test}}))\geq 1-\alpha$ (1)
2. Nonconformity Score Definition and Quantile Calibration
The nonconformity score $s_{t}$ reflects the model’s uncertainty in predicting the true label, defined as:
$s_{t}=1-\hat{f}(X_{t})_{Y_{t}^{*}}$ (2)
Sort the scores of the calibration set in ascending order: $\{s_{1}\leq s_{2}\leq\cdots\leq s_{n}\}$, and Get the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$ quantile from $\{s_{t}\}_{t=1}^{n}$:
$\hat{q}=\inf\left\{q:\frac{|\{t:s_{t}\leq q\}|}{n}\geq\frac{(n+1)(1-\alpha)}{n}\right\}=s_{\frac{\lceil(n+1)(1-\alpha)\rceil}{n}}$ (3)
Key Derivation: Based on the exchangeability assumption, the nonconformity score $s_{\text{test}}$ of the test sample is uniformly distributed among the calibration scores in the sorted sequence. The coverage probability strictly satisfies:
$\mathbb{P}(s_{\text{test}}\leq\hat{q})=\frac{\lceil(n+1)(1-\alpha)\rceil}{n+1}\geq 1-\alpha$ (4)
[…]
3. Prediction Set Construction and Coverage Validation
For the test sample $X_{\text{test}}$, the prediction set includes all classes satisfying the following condition:
$\mathcal{C}(X_{\text{test}})=\left\{y\in[K]:1-\hat{f}(X_{\text{test}})_{y}\leq\hat{q}\right\}$ (5)
Example Operation:
  • 1.
    If the predicted probability for a class $\hat{f}(X_{\text{test}})_{y}\geq 1-\hat{q}$, the class is included in the prediction set.
  • 2.
    The final prediction set size is dynamically adjusted, balancing confidence and classification granularity.
3.1.2 Mathematical Proof Supplement
Theorem (Marginal Coverage Guarantee): Under the data exchangeability assumption, the SCP framework satisfies:
$\mathbb{P}(Y_{\text{test}}^{*}\in\mathcal{C}(X_{\text{test}}))\geq 1-\alpha$ (6)
Proof: Let the combined sequence of calibration and test scores be $s_{(1)}\leq s_{(2)}\leq\cdots\leq s_{(n+1)}$, with the rank $k$ of the test score $s_{\text{test}}$ following a uniform distribution $k\sim\text{Uniform}\{1,\dots,n+1\}$. Select the threshold $\hat{q}=s_{(k^{*})}$, where $k^{*}=\lceil(n+1)(1-\alpha)\rceil$, then:
$\mathbb{P}(s_{\text{test}}\leq\hat{q})=\mathbb{P}(k\leq k^{*})=\frac{k^{*}}{n+1}\geq 1-\alpha$ (7)
3.2 Risk Control in speech emotion recognition task
To adapt the framework of split conformal prediction to the Speech emotion recognition task for user - specified guarantees of task - specific performance, we define a loss function for each calibration data, formulated as
$\ell(C(X_{t}),Y_{t})=\mathbf{1}\{Y_{t}\notin C(X_{t})\}$ (8)
Figure 1: Conformal Risk Control Framework
Conformal Risk Control (CRC) provides statistical guarantees for high - risk classification tasks through the exchangeable data assumption and custom - defined loss functions. Given $N$ calibration data points and 1 test data point (with a total of $K$ classes), CRC first constructs a prediction set for the sample $x_{i}$:
$C_{\beta}(x_{i})=\left\{y:F(x_{i})_{y}\geq 1-\beta\right\}$ (9)
where $F(x_{i})$ is the probability distribution output by the model, and $\beta$ controls the confidence threshold. The task - specific loss is defined as $l_{i}=1-{|C_{\beta}(x_{i})\cap y_{i}^{*}|}$ (if the true label $y_{i}^{*}\notin C_{\beta}(x_{i})$, then $l_{i}=1$), and its expectation must satisfy $\mathbb{E}[l_{\text{test}}]\leq\alpha$. Based on data exchangeability, the expectation of the test loss can be expressed as:
$\mathbb{E}[l_{\text{test}}(\beta)]=\frac{NL_{N}(\beta)+l_{\text{test}}(\beta)}{N+1}\leq\alpha$ (10)
where $L_{N}(\beta)=\frac{1}{N}\sum_{i=1}^{N}l_{i}$ is the average loss of the calibration set. By solving for the optimal threshold
$\hat{\beta}=\inf\left\{\beta:\frac{NL_{N}(\beta)+B}{N+1}\leq\alpha\right\}$ (11)
(where $B$ is the task - related loss upper bound, e.g., $B=1$), $\hat{\beta}$ is determined and applied to the test data to construct $C_{\hat{\beta}}(x_{\text{test}})$, ensuring risk control. This framework flexibly designs loss functions (such as coverage loss $l_{i}=1-$ coverage or sparsity loss $l_{i}=|C_{\beta}(x_{i})|$), transforming abstract task metrics (like false positive rate and prediction set size) into an optimizable form, thereby achieving task - specific metric guarantee.
3.3 Comparison of two frameworks
Conformal Risk Control has a core drawback compared to Split Conformal Prediction in terms of lower computational efficiency, mainly due to its need for multiple traversals of the calibration data to optimize the threshold. Specifically:
The threshold calculation of SCP only requires a single sorting of the non-conformity scores $s_{t}=1-\hat{f}(X_{t})Y^{*}_{t}$ on the calibration set, and taking the quantile $\hat{q}$, with a time complexity of $O(N\log N)$
CRC needs to traverse candidate thresholds $\beta$. For each candidate value, it calculates the average loss on the calibration set $L_{N}(\beta)=\frac{1}{N}\sum_{i=1}^{N}l_{i}$, and then verifies the condition $\frac{NL_{N}(\beta)+B}{N+1}\leq\alpha$. Assuming the number of candidate $\beta$ is $M$, its time complexity is $O(M\cdot N)$ Even with binary search optimization (iteration times $O(\log M)$), it still requires multiple complete traversals of the calibration set, leading to a significant increase in computational load.
This efficiency bottleneck is particularly prominent in large-scale calibration sets ($N\gg 1$) or high-real-time scenarios (such as autonomous driving), while the lightweight single sorting of SCP has more advantages. The flexibility of CRC comes at the cost of efficiency, and one needs to balance the choice according to task requirements.
(a) Training and testing results of Mel-spectrogram features in basic CNN model
(b) Training and testing results of chromaticity features in the basic CNN model
(c) Training and testing results of contrast features in the basic CNN model
Figure 2: Comparative analysis of the performance of different speech features in basic CNN models based on the TESS dataset
4 Experiments
4.1 Experimental Settings
4.1.1 Datasets
We used the Interactive Emotional Binary Motion Capture Database (IEMOCAP) and the Emotional Speech Recognition Dataset (TESS) for model training and evaluation. IEMOCAP contains about 12 hours of audio - visual data, including video, speech, facial motion capture, and text transcription. The data is performed by 10 actors (5 men and 5 women) in 5 sessions, covering emotion categories such as happiness, anger, sadness, frustration, and neutral. TESS is recorded by two actresses (26 and 64 years old, respectively), using the carrier phrase ”Say the word _” to express 200 target words, covering seven emotions: anger, disgust, fear, happiness, surprise, sadness, and neutral. There are a total of 2,800 audio files in WAV format.
(a) ResNet50 training and testing Loss-Acc
(b) MobileNetV3 training and testing Loss-Acc
(c) ShuffleNetV2 training and testing Loss-Acc
(d) SqueezeNet training and testing Loss-Acc
(e) GhostNet training and testing Loss-Acc
Figure 3: Comparison of training and test set Loss-Acc of different neural network models
4.1.2 Base Models
In this study, we used several convolutional neural network (CNN) architectures, including basic CNN, ResNet50, MobileNetV3, ShuffleNetV2, SqueezeNet, and GhostNet. We ensured that the number of parameters of these models was similar: ResNet50 (1,168,955 parameters), MobileNetV3 (1,172,034 parameters), ShuffleNetV2 (1,177,898 parameters), SqueezeNet (1,177,662 parameters), and GhostNet (1,181,503 parameters). These models optimize computational efficiency and performance through different designs. ResNet50 uses residual connections to solve the gradient vanishing problem of deep networks, MobileNetV3 and ShuffleNetV2 use lightweight designs to adapt to resource - constrained environments, and SqueezeNet and GhostNet further optimize the number of model parameters and computational complexity, making them suitable for embedded and mobile device applications.
4.1.3 Methods for Speech Feature Extraction
We compared three speech feature extraction methods: Mel spectrogram, chromaticity feature, and contrast feature. Mel spectrogram is a commonly used feature in speech processing. It maps the spectrum of speech signals to the Mel scale, which simulates the human auditory system’s ability to perceive different frequencies. By converting the speech signal into a spectrum obtained by short - time Fourier transform (STFT) and applying Mel filter bank processing, Mel spectrogram can effectively extract important frequency components in speech. Compared with traditional linear spectrum, Mel spectrogram can better capture the acoustic characteristics of speech. Chromaticity feature converts spectrum information into energy distribution of 12 scales, representing the change of pitch. Through these features, the tonality information in speech can be reflected, which is particularly important for emotion recognition because different emotions may affect the pitch and tonality of speech. Contrast feature reflects the energy difference between different frequency bands in the spectrum and can describe the local structure of the spectrum. It helps capture the detailed changes in speech by calculating the energy difference between each frequency band and the energy of adjacent frequency bands. It can reflect the emotional color and emotional intensity changes of speech.
4.1.4 Evaluation Metrics
We employ the Empirical Coverage Rate (ECR) to evaluate whether we rigorously control the error rates at various user - specified risk levels. Moreover, we leverage the Average Prediction Set Size to assess the uncertainty of the model’s decision - making and the prediction efficiency of the calibrated prediction sets.
4.1.5 Hyper - parameters
We set the split ratio between the calibration set and the test set to 50% in the IEMOCAP dataset and 50% in the TESS dataset. We modified the structure of the models to ensure that each model has a similar number of parameters: ResNet50 (1,168,955 parameters), MobileNetV3 (1,172,034 parameters), ShuffleNetV2 (1,177,898 parameters), SqueezeNet (1,177,662 parameters), and GhostNet (1,181,503 parameters) to evaluate the uncertainty of different models with the same parameter size.
4.2 Determination of Speech Feature Extraction Methods
In the pre - experimental stage, we compared different speech feature extraction methods, including mel - spectrogram features, chromaticity features, and contrast features. We tested the TESS dataset and found that the accuracy of the mel - spectrogram feature on the test set reached 99.82%, significantly better than the chromaticity feature (97.86%) and contrast feature (87.41%). Based on these results, the mel - spectrogram feature showed stronger recognition ability and robustness, so we chose the mel - spectrogram as the optimal feature extraction method and widely used it in the training and evaluation of all models in subsequent experiments.
(a) SCP framework on IEMOCAP
(b) SCP framework on TESS
Figure 4: ECR using SCP framework
(a) Risk Control framework on IEMOCAP
(b) Risk Control framework on TESS
Figure 5: ECR using Risk Control framework
We trained five models (ResNet50, MobileNetV3, ShuffleNetV2, SqueezeNet, and GhostNet), all with a similar number of parameters. During training, mel - spectrogram features were used and trained on the IEMOCAP dataset (60% training set, 40% test set). All models achieved nearly 100% accuracy on the training set, but performed differently on the test set. Specifically, the test set accuracy of ResNet50 was 44.75%, MobileNetV3 was 40.56%, ShuffleNetV2 was 42.38%, SqueezeNet was 37.96%, and GhostNet was 35.06%. These results show that although the models performed well on the training set, they had weak generalization ability on the test set, reflecting the overfitting phenomenon, which provides a basis for the subsequent use of the Conformal Prediction framework for model evaluation.
4.3 Empirical Results of the ECR metric
We used SCP and Risk Control frameworks to evaluate the coverage performance of five models on the IEMOCAP dataset, and applied these models to the TESS dataset for testing. On the IEMOCAP and TESS datasets, both frameworks provided reliable coverage guarantees and showed that the two frameworks had good generalization capabilities across different datasets.
While the theoretical guarantee of conformal prediction is rigorous, there can be minor fluctuations in practice due to finite - sample variability.
4.4 Uncertainty Estimation of the Model through the APSS metric
Dataset Model Risk Level
$0.1$ $0.2$ $0.3$ $0.4$ $0.5$ $0.6$ $0.7$ $0.8$ $0.9$ $1$
IEMOCAP ResNet50 $3.757$ $2.856$ $2.186$ $1.619$ $1.176$ $0.839$ $0.560$ $0.318$ $0.139$ $0.025$
MobileNetV3 $4.158$ $3.132$ $2.440$ $1.847$ $1.369$ $0.982$ $0.645$ $0.380$ $0.155$ $0.066$
ShuffleNetV2 $3.964$ $2.991$ $2.306$ $1.738$ $1.310$ $0.914$ $0.597$ $0.345$ $0.151$ $0.050$
SqueezeNet $4.110$ $3.235$ $2.490$ $1.937$ $1.463$ $1.061$ $0.699$ $0.416$ $0.182$ $0.004$
GhostNet $4.357$ $3.457$ $2.752$ $2.169$ $1.650$ $1.214$ $0.808$ $0.456$ $0.200$ $0.011$
TESS ResNet50 $6.000$ $6.000$ $6.000$ $6.000$ $3.152$ $2.324$ $1.601$ $0.767$ $0.312$ $0.234$
MobileNetV3 $6.000$ $6.000$ $6.000$ $3.832$ $3.346$ $2.877$ $2.333$ $1.590$ $0.619$ $0.100$
ShuffleNetV2 $6.000$ $6.000$ $6.000$ $3.578$ $2.875$ $2.104$ $1.308$ $0.576$ $0.192$ $0.140$
SqueezeNet $6.000$ $6.000$ $6.000$ $5.979$ $3.490$ $2.855$ $2.090$ $0.878$ $0.299$ $0.112$
GhostNet $6.000$ $6.000$ $6.000$ $6.000$ $5.892$ $3.385$ $2.685$ $1.524$ $0.497$ $0.031$
Table 1: APSS under different data sets and risk levels
We found a negative correlation between the average prediction set size (APSS) and the risk level. Specifically, as the risk level increases, the APSS value decreases significantly, indicating that the model’s prediction results at high risk levels are more certain and the size of the prediction set is optimized. In this way, APSS is able to effectively evaluate the uncertainty of the classification model on the test set. Larger APSS values indicate that the model has higher prediction uncertainty for certain categories, while smaller APSS values mean that the model has higher confidence in its predictions. Therefore, APSS, as a measure of model uncertainty, demonstrates its potential as a promising benchmark in evaluating the robustness of classification models and provides a valuable reference for further optimizing models.
4.5 Ablation Studies
4.5.1 Various Split Ratios
Figure 6: Coverage Distribution vs Calibration Set Size ($\alpha=0.2$)
This study systematically explores the impact of different data partitioning strategies on model coverage and verifies the robustness of the speech emotion recognition method based on the conformal prediction framework. In the experimental design, we adopt a dynamic calibration mechanism: the confidence threshold $q$ is calculated through the calibration set data, and its statistical validity is verified under different calibration set - test set partition ratios (10%–90%).
The experimental results show that although the calibration set size changes significantly (from 10% to 90%), this method always ensures the empirical coverage of the test set satisfies the theoretical lower bound $1-\alpha$ ($\alpha=0.2$) by dynamically adjusting the confidence threshold $q$. Theoretical analysis and empirical results jointly prove that this method exhibits strong robustness to data partition ratios. The constructed prediction sets provide statistically rigorous reliability guarantees for speech emotion recognition tasks, effectively addressing the confidence calibration failure problem in traditional methods caused by data distribution shifts.
5 Conclusion
This study addresses the issue of unreliable predictions in traditional neural networks for speech emotion recognition by proposing a solution based on CP and a Risk Control framework. The results indicate that traditional point prediction methods exhibit significant confidence bias due to factors such as data distribution differences, model overfitting, and difficulty in capturing complex emotional features. By using the SCP framework and the risk control framework, this study successfully transforms single-point predictions into set predictions with coverage guarantees, ensuring that, at a specified risk level $\alpha$, the true label is included in the prediction set with at least a probability of $1-\alpha$. Additionally, cross-dataset validation (where a model trained on the IEMOCAP dataset still maintains coverage guarantees on the TESS dataset) demonstrates the robustness of the framework under different data distributions.
We found a significant negative correlation between APSS and risk level, which provides an actionable metric for quantifying model decision uncertainty. Cross-dataset validation results show that even on the TESS dataset, with a large distribution difference, this metric still effectively reflects changes in model confidence. By constructing a dynamic relationship model between the prediction set size and risk level, this study establishes the first uncertainty evaluation standard suitable for speech emotion recognition.
This study adopts a set prediction method based on inconsistency score, and achieves strict coverage guarantee by dynamically adjusting the prediction set threshold. In addition, a more flexible risk control framework is introduced to balance the prediction set size and coverage probability by improving the loss function, thereby optimizing decision efficiency while ensuring statistical reliability. This framework can also be applied to problems with different needs by modifying the definition of the loss function. The study also found that APSS is negatively correlated with the risk level, indicating that when the risk level is high, the model’s prediction set is usually smaller and the model’s prediction uncertainty is lower. Based on this finding, a metric for evaluating the uncertainty of classification models on the test set is proposed.